OpenStudy (anonymous):

Factor x2 - x - 90 = 0

OpenStudy (anonymous):

x^2 - x - 90 = 0 (x - 10)(x + 9) = 0 So, x = 10, x = -9

OpenStudy (anonymous):

wouldnt it be -10?

OpenStudy (anonymous):

and +9

OpenStudy (anonymous):

You would stop at (x-10)(x+9) if they are just looking for the factors of x^2 - x -90

OpenStudy (anonymous):

The solutions to the equation that madibear gave are correct, though, because of the zero product property--if you plug in 10 for x you get: (10 - 10)(10 + 9)=0 (0)(10+9)=0 and anything multiplied by zero = 0, so: 0 = 0

OpenStudy (anonymous):

The same applies for -9, because (x + 9) becomes 0